If n! is divisible by 1532 , the minimum value of n is
Answer by Edwin McCravy(20055)

1532 = 2*766 = 2*2*383

Since 383 is prime, no factorial before 383! contains it as a
factor. But 383! contains 2 as a factor at least twice so the
answer is

383! =

526171930465521393785596080975482340028727379802712622496500876
963285950823045759568621510826669217686959095980120129991997296
111678847138618689553743676623200608630139185868483749406451972
652797372080601186350469827662118721278054876618333025641989470
747253853843684403188032652737687978992980782382630667334670925
071634771933501959569717412491518202250599581857329817250305066
354126336873559016848571122757164611918020749186865176435755672
885471377690987038232530545883783261740691642714422780247302535
241869482325905128700979075906976819859564235316258788734986595
975023130500784811526393017472968767150759365718618086071635135
868396774834248250349890800070786168030132600866093831008299165
773242612121038057040894685301954314240000000000000000000000000
000000000000000000000000000000000000000000000000000000000000000
000000

Edwin